OpenStudy (jessiegonzales):

MEDAL!!! 1. Solve the equation. 3y+20=3+2y

imqwerty (imqwerty):

3y+20 = 3+2y y=3-20 y=-17

OpenStudy (jessiegonzales):

A. -1/17 B. 7 2/3 C. 23 D. -17

OpenStudy (jessiegonzales):

THANKYOU!!!